# Can EDA Tool Feedback Improve Verilog Generation by LLMs?

### v1

@@ 5.2. Mixed-Model Results
Given the relative success, but matching relative expense, of GPT-4o in solving the VerilogEval benchmarking problems, we sought to evaluate if combining a small model with a larger model, such as GPT-4o, could achieve improved results with only minimal impact on cost. Section 5.2 shows the results for ensembling Claude Haiku, GPT-3.5-Turbo, and GPT-4o-Mini respectively with GPT-4o as the final iteration.
@@
Evaluating these results based on model effort, as done with the single-model results, is performed by analyzing the average number of tokens needed to completely generate a benchmark design. Figure 10(a) and Figure 10(b) show these plots.
@@ 5.3. Discussion
RQ2: As the number of candidate responses and the depth of the tree increased we saw a trend higher rates of success with all examined models. The tree search depth, while impactful, seemed to have a smaller effect on the rate of success than the number of candidates did, as evidenced by the rate of success for zero-shot results with many candidates. Increasing the depth, however, resulted in generally fewer tokens needed for the increase as compared to increasing the number of candidates.
@@ 6. Conclusion
Given recent advances in LLM capabilities for both coding and hardware design, it has seemed reasonable to assume that providing a model feedback on its generated code would improve its performance. However, generating this feedback, and ‘explaining’ to the model how and why it is wrong, has typically been done with a human engineer—something costly, nebulous, and potentially slow if the human engineer needs to be able to find and decipher the error on their own before prompting the LLM. As such, systematic studies in this area have been lacking.
In this work, we therefore sought to evaluate how a set of modern, state of the art, general knowledge LLMs would respond to feedback only from EDA tools and testbenches. We sought to discover whether LLMs would truly be able to solve bugs in their own generated designs without the assistance of a knowledgeable human. We further asked: If so, how much effort would it take? How much would it cost? And, are there techniques we can use to make this process better?

### v3

@@ 5.2. Mixed-Model Results
Given the relative success, but matching relative expense, of GPT-4o in solving the VerilogEval benchmarking problems, we sought to evaluate if combining a small model with a larger model, such as GPT-4o, could achieve improved results with only minimal impact on cost. Section 5.2 shows the results for using Claude Haiku, GPT-3.5-Turbo, and GPT-4o-Mini for the complete tree search with GPT-4o as only the final iteration in cases where the maximum depth was reached.
@@
Evaluating these results based on model effort, as with the single-model results, is done by analyzing the average number of tokens needed to completely generate a benchmark design. Figure 12(a) and Figure 12(b) show these plots.
@@ 5.3. Discussion
RQ2: As the number of candidate responses and the depth of the tree increased we saw a trend higher rates of success with all examined models. The tree search depth, while impactful, seemed to have a smaller effect on the rate of success than the number of candidates did, as evidenced by the rate of success for zero-shot results with many candidates. Increasing the depth, however, resulted in generally fewer tokens needed for the increase as compared to increasing the number of candidates. Future work would seek to better explore the reasons for this and what hyperparamters might impact this outcome, as identifying an ideal number of candidates and tree-search depth would potentially help optimize the efficiency of generating function designs even more. Further, an examination of multiple tree-search methodologies, such as MCTS, could shed additional light on the most effective ways to use these models, though such an experiment was considered out of scope for this work.
@@ 6. Conclusion
Given recent advances in LLM capabilities for both coding and hardware design, it has seemed reasonable to assume that providing a model feedback on its generated code would improve its performance. However, generating this feedback, and “explaining” to the model how and why it is wrong, has typically been done with a human engineer—something costly, nebulous, and potentially slow if the human engineer needs to be able to find and decipher the error on their own before prompting the LLM. As such, systematic studies in this area have been lacking.
In this work, we therefore sought to evaluate how a set of modern, state of the art, general knowledge LLMs would respond to feedback only from EDA tools and testbenches. We sought to discover whether LLMs would be able to solve bugs in their own generated designs without the assistance of a knowledgeable human. We further asked: If so, how much effort would it take? How much would it cost? And, are there techniques to improve this process?